OpenStudy (ujjwal):

Electron beam is accelerated from rest through pd 5000 V and impinge normally onto the wall of building with an incident current 25\(\mu\)A. What force is exerted on wall?

OpenStudy (anonymous):

final kinetic energy of electron= charge * pd = e*V= 5000e final momentum=sqrt(2m * 5000e) in time t , the total loss of momentum =m*v*I*t/e force = rate of change of momentum=d/dt (m*v*I*t/e) = mvI/e = sqrt(2m * 5000e) * I/e use